A 23.0 g block of copper whose temperature is 447 K is placed in an insulating box with a 92.4 g block of lead whose temperature is 132 K. (a) What is the equilibrium temperature of the two-block system, in kelvins? (b) What is the change in the internal energy of the two-block system between the initial state and the equilibrium state? (c) What is the change in the entropy of the two-block system? The heat capacities of copper and lead are 386 J/kg
Answer #1

For your convinence

a)23*386*(447-T)/1000=92.4*128*(T-132)/1000

T=267K

b)Change in internal energy=M*C*(Tfinal-Tintial)

=23*386*(267-447)/1000 + 92.4*128*(267-132)/1000

=0

C)Change in entophy is=M*C*Ln(Tfinal/Tintial)

=23*386*Ln(267/447/)/1000 + 92.4*128*Ln(267/132)/1000

=3.756 KJ/K